2x2y=-3
2x-y=5
2x+2(2x-5)=-3
2x+4x-10=-3
6x=7
x=7/6 I understand up to here but then I am confused they got solution 7/6,-8/3 can you please show me step by step please?

Answer by Alan3354(69443)

2x2y=-3
I think you mean 2x+2x = -3
------------
2x-y=5
y = 2x - 5 ********* You said step-by-step, but you left that out.
2x+2(2x-5)=-3 Sub for y
2x+4x-10=-3
6x=7
x=7/6
-----------
I understand up to here but then I am confused they got solution 7/6,-8/3 can you please show me step by step please?
y = 2x - 5
y = 2(7/6) - 5 = 14/6 - 30/6
y = -16/6
y = -8/3